OpenStudy (anonymous):

Last year, a large trucking company delivered 5.5 * 10^5 tons of goods with an average value of $23,000 per ton. What was the total value of the goods delivered? Write the answer in scientific notation. A. 1.265 * 10^9 dollars B. 1.265 * 10^10 dollars C. 1.265 * 10^12 dollars D. 1.265 * 10^11 dollars A, B, C, or D? Please explain

OpenStudy (anonymous):

Tons of goods = 5.5*10^5 Cost of 1 ton = 23,000 Cost of all the goods = 23,000*(5.5*10^5) Which therefore is, 12650000000 and you need to convert it to scientific notation. When you write in this form 1.265*10^? How many numbers do you have after the decimal point? That will be your answer for1.265*10^?